OpenStudy (nuccioreggie):

A set of equations is given below: Equation M: y = 3x + 4 Equation P: y = 3x + 7 Which of the following options is true about the solution to the given set of equations? No solution One solution Two solutions Infinite solutions

OpenStudy (mathewr.):

No solution

OpenStudy (mathewr.):

Medal if helpful

OpenStudy (whovianchick):

y= 3x+4 y=3x+7 Set them equal to each other 3x+4=3x+7 3x=3x+3 0=3 Since 0 doesn't actually equal 3, it's no solution.